Project

General

Profile

Meeting 6-6-2017

Date: 6th June 2017
Time: 15:00 UK time
Attendance: Paul, Colin, Alan, Melissa, Adam,
Apologies:

Agenda and Minutes

Prev Actions:

New Actions:

Tracker Data Review Low Level, MAU

Status of cryostats and Further works, AB, EO CW et al

Proposal that we warm the trackers up over the summer and repeat the recent drying techniques, for longer and perhaps at slightly higher heater temperatures.
There are some interesting left/right asymmetries which it would be good to understand or better still correct.

Agreed to do the pumping over the summer. Colin has two crostats with large bore pipes. We can also use an rga. Will warm up when we can take them to 320 today and then start the pumping as soon as possible. Last time was 320. Need to check the oxford limits are still there. Leak checks to be done at warm.

We can potentially get to 2-4% dead channels using waveguide swaps. cryo 2 still leaks, we need to find and fix this.

Tracker Software Work, PK and AD et al

Work is ongoing. There are a list of actions though, but the main cause for concern atm is that the p value distribution is now peaked at 0 and wrong Since Chris Hunts new calculation. Chris Hunt is going to fix it. We must then do a new release and ensure tests are ready to prevent this happening in future. Kalman needs a lot more tests and documentation.

AoB

Phone Details: International Toll-Free Dial-in numbers

Access code: 1779339517

Dial-in Numbers:
Country Toll-free Toll
  • Bulgaria 008001194419
  • China 8008703563
  • Italy 800149831 0236009716
  • Japan 006633812547
  • Netherlands 08000203285
  • Russia 81080022671012
  • Switzerland 0800920039 0445807737
  • UK 08082380274 02034333570
  • US 631.621.5253 877.914.2048

9412_SciFiReduce_TKU_DigsPerChannel.png (30.3 KB) George, Melissa, 07 June 2017 15:05

9412_SciFiReduce_TKD_DigsPerChannel.png (30 KB) George, Melissa, 07 June 2017 15:05

Deadchannels2017.pdf (449 KB) Overton, Edward, 07 June 2017 15:12